gababb09:

Find the value of the following expression: \[(2^8×~5^{-5}~×19^0)^{-2}×\left(\begin{matrix}5^{-2} \\ 2^{3}\end{matrix}\right)^{4}×2^{28}\] Write your answer in simplified form. Show all of your steps

darkknight:

Start by evaluating some terms, such as 19^0 = 1 and 2^8 = 256 5^-5 = 1/5^5 which equals 1/3125

darkknight:

And then keep simplifying from there. lmk where u are stuck

darkknight:

does that part say (5^-2/2^3)^4???

darkknight:

The format of the problem looks weird, yk where you have 5^-2 and 2^3 all to the power of 4 Does it say 5^-2 divided by 2^3 or...

gababb09: